In a competitive market, firms sell output at a price of ₵ 20. Marginal productivity per hour of the workers is described by the equation MPL = 40 - L. What is the firm’s demand curve for Labour? If the firm can hire Labour from a competitive Labour market at a wage of ₵5 per hour, how many workers should the firm hire?


1
Expert's answer
2021-09-23T17:36:27-0400

a)

Given

Price of output=20

MPL=40-L

Demand curve will be given as

"MPL=\frac{W}{P}\\40-L=\frac{W}{P}"

Where W=wage

P=price of output


b)

Wage=5

"40-L=\frac{W}{P}\\40-L=\frac{5}{20}\\40-0.25=L\\L=39.75"